wolflover1819:

It's for algebra 2 Solve this application using logarithms. How much principal, to the nearest dollar, should you invest at 4% in order to have $3,000 for a vacation to Europe in 3 years?

MusicGeek:

So, the rate of interest would be 4% and the number of years is 3. The formula being used would be A=P(1+r)^t. When you plug the values in, you get 3000=P(1+0.04)^3. Now you just gotta solve for P. 3000=P(1.04)^3. Then you divide both sides by (1.04)^3. 3000/(1,04)^3. Then you get P=$2666.99.